# The non-thermal emission following GW170817 is consistent with a conical radially-stratified outflow with initial Lorentz factor $\lesssim 10$

### v1

@@ Abstract
We show that the non-thermal radio to X-ray emission following the neutron star merger GW 170817 is consistent with synchrotron emission from a collisionless shock driven into the interstellar medium (ISM) by a conical radially stratified outflow observed $\approx 0.25$ rad off-axis, with a power-low mass dependence on momentum, $M(>\gamma\beta)\propto(\gamma\beta)^{-5}$, maximum Lorenz factor $\gamma\approx 10$, opening (half-)angle $\approx 0.2$ rad, and total energy of $\approx 10^{51}$erg. The temporal dependence of the flux during its rising phase is determined by the radial stratification structure, which determines the rate at which outflow energy is deposited in the ISM. This is in contrast with highly relativistic, $\gamma\approx 100$, structured jet models, where the angular jet structure determines the time dependence through the gradual "unveiling" by deceleration of larger angular sections of the jet (which are initially "hidden" by relativistic beaming), typically leading to a predicted flux decline after the peak that is faster than observed. Our model predicts a dependence on the observing angle, which is different than that predicted by highly relativistic jet models. Particularly, similar merger events observed closer to the symmetry axis are predicted to show a similarly extended duration of flux increase with time. Our analysis demonstrates that the data do not require a highly relativistic $\gamma\approx 100$ component, but the presence of such a component with opening angle $\ll 0.2$ rad and energy $\ll 10^{51}$ erg cannot be excluded.
keywords:
(transients:) neutron star mergers – (transients:) black hole-neutron star mergers – gravitational waves – radio continuum: transients – relativistic processes
pubyear: 2024pagerange: The non-thermal emission following GW170817 is consistent with a conical radially-stratified outflow with initial Lorentz factor $\lesssim 10$Data Availability
@@ 1 Introduction
GW170817 provided the first opportunity for a detailed investigation of the off-axis synchrotron emission driven by a collimated relativistic outflow, with a long duration rich set of observations of the non-thermal radio, optical and X-ray emission (Hallinan et al., 2017; Troja et al., 2017; Lyman et al., 2018; Troja et al., 2020; Makhathini et al., 2021) and a measurement of emission centroid motion (Mooley et al., 2018a; Ghirlanda et al., 2019; Mooley et al., 2022).
@@
  • The radio center of light position was measured by the VLBI at 75, 207, and 230 days after the merger (Mooley et al., 2018a; Ghirlanda et al., 2019), complemented by an astrometric measurement of the optical kilonova location eight days after the merger. These measurements indicate a superluminal motion of the center of light implying a $\gamma\gtrsim 5$ outflow (Mooley et al., 2018a, 2022).
@@
The moderate $F_{\nu}\propto t^{0.9}$ flux rise is inconsistent with that predicted for an ultra-relativistic jet with a uniform distribution of energy per solid angle ("top-hat") observed off-axis (Margutti et al., 2018; Mooley et al., 2018b; Gill et al., 2019). Furthermore, the gamma-ray emission associated with GW170817, GRB 170817A, is also inconsistent with such an interpretation (Kasliwal et al., 2017; Granot et al., 2017; Matsumoto et al., 2019). Thus, jets with various angular structures were considered in an attempt to explain the observations (Troja et al., 2019; Fong et al., 2019; Lamb et al., 2019; Hajela et al., 2019; Wu & MacFadyen, 2019; Ryan et al., 2020; Takahashi & Ioka, 2020).
@@
In the highly relativistic "structured jet" models, the jet properties (e.g. energy per unit solid angle) vary as a function of the angle with respect to the jet axis. Most papers fitting structured jet models to the obsreved non-thermal light curve of GW170817 use semi-analytic codes, such as "afterglowpy", for the emission predicted for given model parameters, combined with an MCMC parameter estimation pipeline (Hajela et al., 2019; Lamb et al., 2019; Ghirlanda et al., 2019; Wu & MacFadyen, 2019; Troja et al., 2019; Ryan et al., 2020; Troja et al., 2020; Makhathini et al., 2021; Ryan et al., 2023; McDowell & MacFadyen, 2023; Palmese et al., 2024; Pellouin & Daigne, 2024; Morsony et al., 2024; Wang et al., 2024; Nedora et al., 2024). These models typically cannot account simultaneously for both the light curve and the centroid motion, and they typically do not include an accurate treatment of the jet sideways expansion, which considerably modifies the predictions for the light curve decline phase. Furthermore, the jet energy inferred in some of these models is typically $\sim 10^{52}$ erg, a relatively high value that corresponds to an isotropic equivalent energy $>10^{54}$ erg, which is inconsistent with short GRB observations (D’Avanzo et al., 2014).
@@
A more numeric approach to fitting structured jet models to the data was adopted in several papers, based on using numeric 2D hydrodynamic calculations (Mooley et al., 2018a; Gill et al., 2019; Gottlieb et al., 2019), approximate 1D hydrodynamic calculations (Mooley et al., 2022), and numerically calibrated (to ultra-relativistic structured jet calculations) analytic models (Govreen-Segal & Nakar, 2023). These analyses typically find a jet opening angle of $\theta_{j}\sim 3^{\circ}-6^{\circ}$ and a viewing angle of $\theta_{\text{v}}\sim 15^{\circ}-30^{\circ}$ with respect to the jet axis. The decline of the flux after its peak time is typically predicted in these models to be faster than observed up to $\sim 1000$ days. This is due to the fact that the phase of increasing observed flux extends up to the time when deceleration unveils the full angular extent of the jet to the observer, and is then followed by a rapid flux decline that gradually shallows to an asymptotic $F_{\nu}\propto t^{-p}$ behavior, while the observed decline is shallower than $t^{-p}$ (with $p=2.17$).
@@
In this paper, we show that the non-thermal emission is consistent with synchrotron emission from a collisionless shock driven into the ISM by a conical radially stratified outflow observed off-axis, with a power-low mass dependence on momentum, $M(>\gamma\beta)\propto(\gamma\beta)^{-5}$ and $\gamma\lesssim 10$. In this model, the peak in observed flux is obtained when the reverse shock, which is driven into the ejecta and decelerates it due to the interaction with the ISM, crosses through the ejecta (for highly relativistic, $\gamma\approx 100$, outflow the reverse shock crosses the ejecta and all the energy is transferred to the ISM at smaller radii, $r\propto\gamma^{-2/3}$, corresponding to much earlier observed times preceding the first radio detection). A shallow decline past peak is obtained for radially stratified outflows in cases where the reverse shock crosses through the ejecta before the deceleration reveals the full outflow extent to an off-axis observer.
@@
The paper is organized as follows. In § 2, we present a simple semi-analytic method to calculate the synchrotron emission from a conical stratified outflow. 2D numeric calculations are presented in § 3. The accuracy of our semi-analytic calculations is determined in § 4 through a comparison to the numeric results. In all calculations, we assume that electrons are accelerated by the collisionless shock to power-law energy distribution, $dn_{e}/d\gamma_{e}\propto\gamma_{e}^{-p}$ where $\gamma_{e}$ is the electron Lorenz factor (in the plasma rest frame), and that fractions $\varepsilon_{e}$ and $\varepsilon_{B}$ of the post-shock thermal energy are carried by non-thermal electrons and magnetic fields, respectively. In § 5 we show, using both semi-analytic and numeric calculations, that the stratified outflow model may account for the non-thermal emission following GW170817. Our conclusions are summarized in § 6.
@@ 2.1 Initial ejecta structure
$M(>\gamma\beta)=M_{c}\left(\frac{\gamma\beta}{\gamma_{c}\beta_{c}}\right)^{-k}\text{for}\quad\gamma_{c}\beta_{c}<\gamma\beta.$ (1)
@@ 3.1 Hydrodynamics
which is in very good agreement with the equation of state of a mildly relativistic fluid provided by Synge (1957). The $1.1$ factor is introduced for better agreement with the exact solution in the range of $94<e/\rho^{\prime}c^{2}<$ 9 (corresponding to $105<\gamma<$ 10, Sadeh, 2024a).
Our simulations were performed in a 2D cylindrical coordinate system $(r,z)$ assuming axial symmetry, and applying reflective boundary conditions at both the symmetry axis ($r=0$) and the symmetry plane ($z=0$). The computational domain boundaries were set at $r=10^{19}$cm and $z=10^{19}$cm. The initial grid prioritizes resolving the ejecta by placing a higher concentration of cells within it, with a coarser spacing outside. Typically, the simulation begins with $500\times 500$ cells, with most cells within the ejecta. AMR dynamically adjusts the resolution throughout the simulation, optimizing cell distribution along both the $r$ and $z$ axes to effectively capture regions of significant variation in pressure, density, or Lorentz factor. We systematically refined the initial spatial grid and temporal steps to assess convergence. We also iteratively increased the number of times the regridding scheme is allowed to multiply the number of cells within a time step. This process was continued until no significant changes were observed in the hydrodynamics and emergent flux.
The initial ejecta structure was chosen as described in § 2.1. The initial radius of a mass shell $M(>\gamma\beta)$ was determined as $r_{0}=\beta ct_{0}$, with $t_{0}\approx 10^{6}$s. Behind the ejecta, between $r=\beta_{c}ct_{0}$ and $r=0.99\beta_{c}ct_{0}$, we chose a smooth power-law interpolation between $\beta_{c}$, $\rho_{c}$ (the density of slowest part the ejecta) and $\beta=0,\rho=10^{-3}\rho_{\text{ISM}}$ to avoid numeric instabilities due to discontinuities in the density and velocity. The initial pressure in all of the cells is set to $P=10^{-10}\rho_{\text{ISM}}c^{2}$ where $\rho_{\text{ISM}}$ is the ISM mass density.
@@ 3.2 Radiation
In Fig. 2, we compare the light curve (at frequencies above the self-absorption and peak frequency and below the cooling frequency) and the time-dependent light centroid location obtained semi-analytically with the results of the 2D numerical calculation. Recall that the semi-analytic description is valid up to time at which the reverse shock crosses the ejecta. The agreement is within $10$’s of percent for relevant values of $\{k,\gamma_{c},n,M_{c},\varepsilon_{e},\varepsilon_{B},p,\theta_{\text{open}}\}$.
Figure 2: A comparison between the time-dependent flux (left panel) and the time-dependent light centroid location (right panel) obtained semi-analytically (solid lines) and numerically (dashed lines), for various viewing angles and two sets of parameter values $\{k,\gamma_{c},n,M_{c},\varepsilon_{e},\varepsilon_{B},p,\theta_{\text{open}}\}=$ A: $\{5,3,10^{-2}\text{cm}^{-3},10^{-1}M_{\odot},10^{-1},10^{-2},2.2,30^{\circ}\}$, B: $\{5,3,10^{-3}\text{cm}^{-3},10^{-2}M_{\odot},10^{-1},10^{-2},2.2,20^{\circ}\}$. The flux is shown for a distance of $100$Mpc and a frequency of 6 GHz.
@@ 5 GW170817
In Sadeh (2024b), we show that the absence of breaks from the non-thermal power-law spectrum of GW170817 implies that (i) the Lorentz factor of the plasma emitting the observed radiation is $\gamma>2.6$ at $t\sim 10$ days and $\gamma<12$ at $t>16$ days, and (ii) $n\cdot\varepsilon_{B}\lesssim 3\times 10^{-7}$cm-3. Considering these constraints, we used our semi-analytic method to find parameter values that yield model predictions that are in reasonable agreement with observations. We then used the 2D numeric code to calculate more accurately the model predictions, and to extend them to times past the validity time of the semi-analytic approximation. We do not attempt to obtain a "best-fit" set of parameter values, as the real outflow structure is likely more complicated than the simplified structure that we are using for our model calculations, which also includes simplifying assumptions regarding the microphysics of particle acceleration and emission of radiation (e.g. fixed electron and magnetic field energy fractions, independent of shock parameters) We found that the following choice of parameters yields a good fit to the data: $\{k,\gamma_{c},n,M_{c},\varepsilon_{e},\varepsilon_{B},p,\theta_{\text{open}},\theta_{\text{v}}\}=\{4.8,6,5\times 10^{-4}\text{cm}^{-3},9\times 10^{-3}M_{\odot},0.1,2\times 10^{-5},2.17,10^{\circ},16^{\circ}\}$ (Makhathini et al., 2021; Mooley et al., 2018b, 2022), as shown in Figs. 3-4. In the numeric calculation, the ejecta is truncated beyond $\gamma=12$.
Figure 3: The non-thermal radio to X-ray flux (circles, Makhathini et al., 2021) of the electromagnetic counterpart of GW170817, compared to the results of a 2D numeric calculation of the emission driven by a stratified conical ejecta with $\{k,\gamma_{c},n,M_{c},\varepsilon_{e},\varepsilon_{B},p,\theta_{\text{open}},\theta_{\text{v}}\}=\{4.8,6,5\times 10^{-4}\text{cm}^{-3},9\times 10^{-3}M_{\odot},0.1,2\times 10^{-5},2.17,10^{\circ},16^{\circ}\}$. The model ejecta is truncated beyond $\gamma=12$.
Figure 4: The measured motion of the light centroid (circles, Mooley et al., 2018b, 2022) compared to the results of a 2D numeric calculation of the emission driven by a stratified conical ejecta with $\{k,\gamma_{c},n,M_{c},\varepsilon_{e},\varepsilon_{B},p,\theta_{\text{open}},\theta_{\text{v}}\}=\{4.8,6,5\times 10^{-4}\text{cm}^{-3},9\times 10^{-3}M_{\odot},0.1,2\times 10^{-5},2.17,10^{\circ},16^{\circ}\}$. The model ejecta is truncated beyond $\gamma=12$.

### v2

@@ Abstract
We show that the non-thermal radio to X-ray emission following the neutron star merger GW170817 is consistent with synchrotron emission from a collisionless shock driven into the interstellar medium (ISM) by a conical radially stratified outflow observed $\approx 0.25$ rad off-axis, with a power-low mass dependence on momentum, $M(>\gamma\beta)\propto(\gamma\beta)^{-4}$, maximum Lorenz factor $\gamma=10$, opening (half-)angle $\approx 0.15$ rad, and total energy of $\approx 5\times 10^{50}$erg. The temporal dependence of the flux during its rising phase is determined by the radial stratification structure, which determines the rate at which outflow energy is deposited in the ISM. This is in contrast with highly relativistic, $\gamma\approx 100$, structured jet models, where the angular jet structure determines the time dependence through the gradual "unveiling" by deceleration of larger angular sections of the jet (which are initially "hidden" by relativistic beaming), typically leading to a predicted flux decline after the peak that is faster than observed. Our model predicts a dependence on the observing angle, which is different than that predicted by highly relativistic jet models. Particularly, similar merger events observed closer to the symmetry axis are predicted to show a similarly extended duration of flux increase with time. Our analysis demonstrates that the data do not require a highly relativistic $\gamma\approx 100$ component, but the presence of such a component with opening angle $\ll 0.15$ rad and energy $\ll 5\times 10^{50}$ erg cannot be excluded.
X-ray transient sources(1852)– Radio transient sources(2008)– Gravitational wave sources(677) – Neutron stars(1108)– Relativistic fluid dynamics(1389)
@@ 1 Introduction
GW170817 provided the first opportunity for a detailed investigation of the off-axis synchrotron emission driven by a collimated relativistic outflow, with a long duration rich set of observations of the non-thermal radio, optical and X-ray emission (Hallinan et al., 2017; Troja et al., 2017; Lyman et al., 2018; Troja et al., 2020; Makhathini et al., 2021) and measurement of emission centroid motion (Mooley et al., 2018a; Ghirlanda et al., 2019; Mooley et al., 2022).
@@
  • The radio center of light position was measured by the VLBI at 75, 207, and 230 days after the merger (Mooley et al., 2018a; Ghirlanda et al., 2019), complemented by an astrometric measurement of the optical kilonova location eight days after the merger. These measurements indicate a superluminal motion of the flux centroid, implying a $\gamma\gtrsim 5$ outflow (Mooley et al., 2018a, 2022).
@@
The moderate $F_{\nu}\propto t^{0.9}$ flux rise is inconsistent with that predicted for an ultra-relativistic jet with a uniform distribution of energy per solid angle ("top-hat") observed off-axis (Margutti et al., 2018; Mooley et al., 2018b; Gill et al., 2019). Furthermore, the gamma-ray emission associated with GW170817, GRB 170817A, is also inconsistent with such an interpretation (Kasliwal et al., 2017; Granot et al., 2017; Matsumoto et al., 2019a). Thus, jets with various angular structures were considered in an attempt to explain the observations (Troja et al., 2019; Fong et al., 2019; Lamb et al., 2019; Hajela et al., 2019; Wu & MacFadyen, 2019; Ryan et al., 2020; Takahashi & Ioka, 2020). We note that wide-angle outflow models were initially proposed to account for the observations (Mooley et al., 2018b; D’Avanzo et al., 2018; Hotokezaka et al., 2018; Nakar & Piran, 2018; Troja et al., 2018; Gill & Granot, 2018), but were set aside in favor of the ’structured jet’ models following the detection of the radio centroid motion (Mooley et al., 2018a), since wide-angle outflow models did not predict the observed fast motion. Furthermore, the post-peak decline predicted by these models is shallower than observed (Makhathini et al., 2021).
@@
In the highly relativistic "structured jet" models, the jet properties (e.g., energy per unit solid angle) vary as a function of the angle with respect to the jet axis. Most papers fitting structured jet models to the observed non-thermal light curve of GW170817 use semi-analytic codes and analytic estimations, such as "afterglowpy", for the emission predicted for given model parameters, combined with an MCMC parameter estimation pipeline (Hajela et al., 2019; Lamb et al., 2019; Ghirlanda et al., 2019; Wu & MacFadyen, 2019; Troja et al., 2019; Ryan et al., 2020; Troja et al., 2020; Beniamini et al., 2020; Makhathini et al., 2021; Beniamini et al., 2022; Ryan et al., 2023; McDowell & MacFadyen, 2023; Palmese et al., 2024; Pellouin & Daigne, 2024; Morsony et al., 2024; Wang et al., 2024; Nedora et al., 2024). Incorporating an accurate treatment of the jet sideways expansion in semi-analytic methods is highly challenging, and it can considerably modify the predictions for the light curve decline phase. Additionally, these models encounter difficulties in simultaneously reproducing both the light curve and centroid motion without an additional late-time component. Furthermore, the jet energy inferred in some of these models is typically $\sim 10^{52}$ erg, a relatively high value that corresponds to isotropic equivalent energy $\gtrsim 10^{54}$ erg (Ryan et al., 2023; Wang et al., 2024; Pellouin & Daigne, 2024; Nedora et al., 2024), which is inconsistent with short GRB observations (D’Avanzo et al., 2014).
@@
A more numeric approach to fitting structured jet models to the data was adopted in several papers, based on using numeric 2D hydrodynamic calculations (Mooley et al., 2018a; Gill et al., 2019; Gottlieb et al., 2019), approximate 1D hydrodynamic calculations (Mooley et al., 2022), and numerically calibrated (to ultra-relativistic structured jet calculations) analytic models (Govreen-Segal & Nakar, 2023). These analyses typically find a jet opening angle of $\theta_{j}\sim 3^{\circ}-6^{\circ}$ and a viewing angle of $\theta_{\text{v}}\sim 15^{\circ}-30^{\circ}$ with respect to the jet axis. The decline of the flux after its peak time is typically predicted in these models to be faster than observed up to $\sim 1000$ days. This is due to the fact that the phase of increasing observed flux extends up to the time when deceleration unveils the full angular extent of the jet to the observer and is then followed by a rapid flux decline that gradually shallows to an asymptotic $F_{\nu}\propto t^{-p}$ behavior, while the observed decline is shallower than $t^{-p}$ (with $p=2.17$).
A note is in place here regarding the $\gamma$-ray flash associated with GRB 170817A (Goldstein et al., 2017; Savchenko et al., 2017), which carries an energy that is smaller by $\sim$4 orders of magnitude compared to that of typical short GRBs (sGRB). The low energy is unlikely due to relativistic beaming suppression, i.e., due to observing a highly relativistic $\Gamma\sim 100$ jet that produces a typical sGRB along its axis, at a large off-axis angle, $\theta-\theta_{j}>1/\Gamma$ where $\theta_{j}$ is the jet opening angle. This is due mainly to the fact that, in this case, the characteristic photon energy is also reduced by a large factor (Granot et al., 2017; Kasliwal et al., 2017; Matsumoto et al., 2019b). The mechanism producing the $\gamma$-ray flash is not known (different explanations suggested, e.g., Murguia-Berthier et al., 2017; Ioka & Nakamura, 2018; Nakar et al., 2018; Gottlieb et al., 2018; Lundman & Beloborodov, 2021; Beloborodov et al., 2020), and it is thus difficult to derive constraints on the outflow based on its properties. Assuming that the photon spectrum follows a "Comptonized" spectrum up to $\sim 1$ MeV, would imply a lower limit of a few to the Lorentz factor of the plasma emitting the radiation (Matsumoto et al., 2019a, b).
@@
In this paper, we show that the non-thermal emission is consistent with synchrotron emission from a collisionless shock driven into the ISM by a conical radially stratified outflow observed off-axis (at an angle comparable to the flow opening angle), with a power-low mass dependence on momentum, $M(>\gamma\beta)\propto(\gamma\beta)^{-4}$ and $\gamma\lesssim 10$. In this model, the peak in observed flux is obtained when the reverse shock, which is driven into the ejecta and decelerates it due to the interaction with the ISM, crosses through the ejecta (for highly relativistic, $\gamma\approx 100$, outflow the reverse shock crosses the ejecta and all the energy is transferred to the ISM at smaller radii, $r\propto\gamma^{-2/3}$, corresponding to much earlier observed times preceding the first radio detection). A shallow decline post-peak is obtained for radially stratified outflows in cases where the reverse shock crosses through the ejecta before the deceleration reveals the full outflow extent to an off-axis observer.
We note that shock breakout through a wide-angle stratified outflow, $M(>\gamma\beta)\propto(\gamma\beta)^{-4}$, was suggested as a source of the $\gamma$-ray flash (Lundman & Beloborodov, 2021; Beloborodov et al., 2020). While the structure they consider is similar to that considered in this paper, the outflow mass required in their model to account for the $\gamma$-ray emission, $10^{-7}M_{\odot}$ at $\gamma>5$, is smaller by $2-3$ orders of magnitude compared to that we infer from the radio emission.
@@
The paper is organized as follows. In § 2, we present a simple semi-analytic method to calculate the synchrotron emission from a conical stratified outflow. 2D numeric calculations are presented in § 3. The accuracy of our semi-analytic calculations is determined in § 4 through a comparison to the numeric results. In all calculations, we assume that electrons are accelerated by the collisionless shock to power-law energy distribution, $dn_{e}/d\gamma_{e}\propto\gamma_{e}^{-p}$ where $\gamma_{e}$ is the electron Lorenz factor (in the plasma rest frame), and that fractions $\varepsilon_{e}$ and $\varepsilon_{B}$ of the post-shock thermal energy are carried by non-thermal electrons and magnetic fields, respectively. In § 5, we show, using both semi-analytic and numeric calculations, that the stratified outflow model may account for the non-thermal emission following GW170817. Our conclusions are summarized in § 6.
@@ 2.1 Initial ejecta structure
$M(>\gamma\beta)=M_{c}\left(\frac{\gamma\beta}{\gamma_{c}\beta_{c}}\right)^{-k}\text{for}\quad\gamma_{c}\beta_{c}<\gamma\beta,$ (1)
where $M_{c}$ is the equivalent isotropic mass of the conical ejecta.
@@ 3.1 Hydrodynamics
which is in very good agreement with the equation of state of a mildly relativistic fluid provided by Synge & Morse (1958). The $1.1$ factor is introduced for better agreement with the exact solution in the range of $94<e/\rho^{\prime}c^{2}<$ 9 (corresponding to $105<\gamma<$ 10, Sadeh, 2024a).
Our simulations were performed in a 2D cylindrical coordinate system $(r,z)$ assuming axial symmetry and applying reflective boundary conditions at both the symmetry axis ($r=0$) and the symmetry plane ($z=0$). The computational domain boundaries were set at $r=10^{19}$cm and $z=10^{19}$cm. The initial grid prioritizes resolving the ejecta by placing a higher concentration of cells within it, with a coarser spacing outside. Typically, the simulation begins with $500\times 500$ cells, with most cells within the ejecta. AMR dynamically adjusts the resolution throughout the simulation, optimizing cell distribution along both the $r$ and $z$ axes to effectively capture regions of significant variation in pressure, density, or Lorentz factor. We systematically refined the initial spatial grid and temporal steps to assess convergence. We also iteratively increased the number of times the regridding scheme is allowed to multiply the number of cells within a time step. This process was continued until no significant changes were observed in the hydrodynamics and emergent flux.
The initial ejecta structure was chosen as described in § 2.1. The initial radius of a mass shell $M(>\gamma\beta)$ was determined as $r_{0}=\beta ct_{0}$, with $t_{0}=10^{6}-10^{7}$s. Behind the ejecta, between $r=\beta_{c}ct_{0}$ and $r=0.99\beta_{c}ct_{0}$, we chose a smooth power-law interpolation between $\beta_{c}$, $\rho_{c}$ (the density of slowest part the ejecta) and $\beta=0,\rho=10^{-3}\rho_{\text{ISM}}$ to avoid numeric instabilities due to discontinuities in the density and velocity. The initial pressure in all of the cells is set to $P=10^{-10}\rho_{\text{ISM}}c^{2}$ where $\rho_{\text{ISM}}$ is the ISM mass density.
@@ 3.2 Radiation
In Fig. 2, we compare the light curve (at frequencies above the self-absorption and peak frequency and below the cooling frequency), and the time-dependent light centroid location obtained semi-analytically with the results of the 2D numerical calculation. Recall that the semi-analytic description is valid up to the time at which the reverse shock crosses the ejecta, as such, we can compare the two only prior to the peak time scale. Unfortunately, there is no simple analytic expression for the temporal dependence of the rise and decline, which are highly sensitive to the velocity profile, the opening, and the viewing angle. The agreement is within $10$’s of percent for relevant values of $\{k,\gamma_{c},n,M_{c},\varepsilon_{e},\varepsilon_{B},p,\theta_{\text{open}}\}$.
Figure 2: A comparison between the time-dependent flux (left panel) and the time-dependent light centroid location (right panel) obtained semi-analytically (solid lines) and numerically (dashed lines), for various viewing angles and two sets of parameter values $\{k,\gamma_{c},n,M_{c},\varepsilon_{e},\varepsilon_{B},p,\theta_{\text{open}}\}=$ A: $\{5,3,10^{-2}\text{cm}^{-3},5.5\times 10^{-4}M_{\odot},10^{-1},10^{-2},2.2,30^{\circ}\}$, B: $\{5,3,10^{-3}\text{cm}^{-3},5.5\times 10^{-5}M_{\odot},10^{-1},10^{-2},2.2,20^{\circ}\}$. Notice that $\theta_{\text{open}}$ represents the initial opening angle of the conical ejecta while $\theta_{V}$ is the viewing angle of the observer. Since the semi-analytic description is valid up to the time at which the reverse shock crosses the ejecta, we compare the two only prior to the peak time scale. The flux is shown for a distance of $100$Mpc and a frequency of 6 GHz.
@@ 5 GW170817
In Sadeh (2024b), we show that the absence of breaks from the non-thermal power-law spectrum of GW170817 implies that (i) the Lorentz factor of the plasma emitting the observed radiation is $\gamma>2.6$ at $t\sim 10$ days and $\gamma<12$ at $t>16$ days, and (ii) $n\cdot\varepsilon_{B}\lesssim 3\times 10^{-7}$cm-3. Considering these constraints, we used our semi-analytic method to find parameter values that yield model predictions that are in reasonable agreement with observations. We then used the 2D numeric code to calculate the model predictions more accurately and extend them to times past the validity time of the semi-analytic approximation. We do not attempt to obtain a "best-fit" set of parameter values, as the actual outflow structure is likely more complicated than the simplified structure that we are using for our model calculations, which also includes simplifying assumptions regarding the microphysics of particle acceleration and emission of radiation (e.g., fixed electron and magnetic field energy fractions, independent of shock parameters). We found that the following choice of parameters yields a good fit to the data: $\{k,\gamma_{c},n,M_{c},\varepsilon_{e},\varepsilon_{B},p,\theta_{\text{open}},\theta_{\text{v}}\}=\{4,5.6,5\times 10^{-4}\text{cm}^{-3},5\times 10^{-3}M_{\odot},0.1,5\times 10^{-5},2.17,8^{\circ},15^{\circ}\}$ (Makhathini et al., 2021; Mooley et al., 2018b, 2022), as shown in Figs. 3-5. In the numeric calculation, the ejecta is truncated beyond $\gamma=10$.
Figure 3: The non-thermal radio to X-ray flux (circles, Makhathini et al., 2021) of the electromagnetic counterpart of GW170817, compared to the results of a 2D numeric calculation of the emission driven by a stratified conical ejecta with $\{k,\gamma_{c},n,M_{c},\varepsilon_{e},\varepsilon_{B},p,\theta_{\text{open}},\theta_{\text{v}}\}=\{4,5.6,5\times 10^{-4}\text{cm}^{-3},5\times 10^{-3}M_{\odot},0.1,5\times 10^{-5},2.17,8^{\circ},15^{\circ}\}$. The model ejecta is truncated beyond $\gamma=10$. The flux is shown for a distance of $41$Mpc.
The measured motion of the light centroid (circles, Mooley et al., 2018b, 2022) compared to the results of a 2D numeric calculation of the emission driven by a stratified conical ejecta with $\{k,\gamma_{c},n,M_{c},\varepsilon_{e},\varepsilon_{B},p,\theta_{\text{open}},\theta_{\text{v}}\}=\{4,5.6,5\times 10^{-4}\text{cm}^{-3},5\times 10^{-3}M_{\odot},0.1,5\times 10^{-5},2.17,8^{\circ},15^{\circ}\}$. The model ejecta is truncated beyond $\gamma=10$. The centroid motion is shown for a distance of $41$Mpc.